Joanna deposited $21,000 in 2 banks. One bank gives 2.5% interest the other gives 3.25% Annual interest. How much did she deposit in each bank if she receives a total of $615 annual interest?
Answer by Boreal(15235)

x=2.5% bank
21000-x=3.25%
.025x+(21000-x)*(0.0325)=615
.025x+682.50-0.0325x=615
-0.0075x=-67.50
x=9000 in bank 1 @ 0.025 or $225 interest
21000-x=12000 in bank 2 @ 0.0325 or $390 interest
adds to $615